OpenStudy (anonymous):

using integration by substitution,what is the integral of 2x cos(x^2-1)dx

OpenStudy (anonymous):

Make x^2 - 1 = u then du = 2xdx, and we are left with\[\int\limits \cos(u) du\]

OpenStudy (anonymous):